OpenStudy (anonymous):

Given f(x)=x^3-4x^2-2. State the intervals on which f is increasing and on which f is decreasing.

OpenStudy (amistre64):

derive and set to 0

OpenStudy (amistre64):

\[f' = 3x^2 -8x =0\] \[x(3x-8)=0\] \(x=0,8/3 \) right?

OpenStudy (amistre64):

then maybe this; but we gotta check to make sure that x=0 aint an inflectionin disguise <.........0...........8/3...........> - + + - - + ------------------------ + - +

OpenStudy (anonymous):

Wait, where are you getting the equation from?

OpenStudy (amistre64):

its called a derivative; the first derivative tells you how the function is moving.... its slope at any given point

OpenStudy (amistre64):

when the derivative =0 that means we are bending back on itself and changing direction ..usually. but, an inflection in a graph can also show a 0 in the 1st derivative

OpenStudy (amistre64):

but; 6x-8 aint =0 at x=0 so its a bender

OpenStudy (amistre64):

i gues the long way would be to factor the equation if we can

OpenStudy (amistre64):

i just assumed by the nature of the question that it was a derivative type question lol

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x^3-4x^2-2 might help you see whats happenin with it

OpenStudy (anonymous):

I'm not up to calc yet :/ sorry...

OpenStudy (anonymous):

Helppppp....